A Complete Bibliography of Publications in Canadian Journal of Mathematics = Journal canadien de

mathématiques for the decade 1940–1949

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.04

Title word cross-reference

abstract [33]. algebraic [5]. Angular [24]. Applications [30]. arbitrary [8]. associated [18]. Axiomatic [31].

between [27]. bounded [11, 14]. Boundedness [16].

Cayleyan [4]. certain [9]. characteristic [35]. combinatorial [30]. Completely [13]. Congruence [27]. conjugate [7]. continuous [12]. convex [7]. courbure [2]. critical [8]. curvature [24, 2].

degree [33, 11]. density [25]. differential [18]. dimensional [10]. Direct [28].

Discrete [3]. disjoint [15].

eigenfunctions [34, 21]. Elementary [5]. Epstein [29]. equation [18].

Factorization [34]. fields [19, 35]. finite [9, 17, 10]. finite-dimensional [10]. force [19]. force-fields [19]. forms [32]. formulae [30]. four [6]. four-space [6]. Fréchet [14]. free [17]. French [2]. function [16, 18]. function-lattices [16]. Functionals [14]. functions [7, 29].

General [20]. generalization [29]. generalizations [30]. Generalized [10, 4]. given [33]. Graphs [33]. Green [18]. group [33, 15]. Groups [11, 17].

Hermann [29].

indecomposable [13]. index [17]. infinite [31]. integers [25]. integral [24, 3]. integrals [12]. irreducible [15]. isotherm [2]. isotherme [2].

ladders [34]. Laplace [21]. Laplace-operator [21]. Latimer [26]. lattices [8, 16]. Liouville [35]. Lorentz [3].

MacDuffee [26]. manifolds [21]. matrices [6]. matrix [27]. Mean [12, 2]. Mean-continuous [12]. measure [24]. mechanical [5]. methods [28]. modules [13]. motion [20, 22]. moyenne [2].

Newtonian [19]. nonexistence [9]. Note [19, 4, 29].

operator [21, 4]. order [18]. Orthogonal [6].

particles [20]. planes [9]. point [8]. positive [35]. power [1]. powers [27]. problem [5]. product [1, 15]. projective [9]. properties [21, 16].

quadratic [32]. quantum [5].

rank [31]. reducible [25]. relations [27]. Relativity [20]. Representation [32]. representations [11, 15]. Riemannian [21].

second [18]. second-order [18]. series [1]. sets [8, 31]. Some [21]. space [6, 3]. space-time [3]. spaces [10]. structure [10]. Subgroups [17]. summability [28]. supplementary [29]. surface [23]. surfaces [2]. symmetric [15]. symmetry [5].

theorem [35, 30, 26]. theorems [28]. Theory [20]. three [33, 22]. time [3]. traces [27]. transformations [3]. treatment [31, 5]. two [1].

uniqueness [18]. upon [4].

variation [14]. vector [10]. vortices [22].

waves [23]. Weyl [29]. Wilson [30].

zeta [29].

References

Davenport:1949:PTP

[1] H. Davenport and G. Pólya. On the product of two power series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(1):1–5, ???? 1949. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lalan:1949:SCM

[2] Victor Lalan. Sur les surfaces à courbure moyenne isotherme. (French) [On surfaces with isotherm mean curvature]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1 (??):6–28, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schild:1949:DST

[3] Alfred Schild. Discrete space-time and integral Lorentz transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):29–47, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Turnbull:1949:NUG

[4] H. W. Turnbull. Note upon the generalized Cayleyan operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):48–56, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weyl:1949:EAT

[5] Hermann Weyl. Elementary algebraic treatment of the quantum mechanical symmetry problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):57–68, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

MacDuffee:1949:OMF

[6] C. C. MacDuffee. Orthogonal matrices in four-space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(1):69–72, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fenchel:1949:CCF

[7] W. Fenchel. On conjugate convex functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):73–77, ???? 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mahler:1949:CLA

[8] K. Mahler. On the critical lattices of arbitrary point sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):78–87, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruck:1949:NCF

[9] R. H. Bruck and H. J. Ryser. The nonexistence of certain finite projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):88–93, ???? 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Menger:1949:GVS

[10] Karl Menger. Generalized vector spaces. I. The structure of finite-dimensional spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1 (??):94–104, ???? 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaplansky:1949:GRB

[11] Irving Kaplansky. Groups with representations of bounded degree. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):105– 112, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1949:MCI

[12] H. W. Ellis. Mean-continuous integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1 (??):113–124, ???? 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snapper:1949:CIM

[13] Ernst Snapper. Completely indecomposable modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):125–152, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morse:1949:FBF

[14] Marston Morse and William Transue. Functionals of bounded Fréchet variation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1 (??):153–165, ???? 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1949:DPI

[15] G. de B. Robinson. On the disjoint product of irreducible representations of the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):166–175, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stone:1949:BPF

[16] M. H. Stone. Boundedness properties in function-lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):176–186, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hall:1949:SFI

[17] Marshall Hall, Jr. Subgroups of finite index in free groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):187–190, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Titchmarsh:1949:UGF

[18] E. C. Titchmarsh. On the uniqueness of the Green’s function associated with a second-order differential equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):191– 198, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Birkhoff:1949:NNF

[19] Garrett Birkhoff and Lindley Burton.

Note on Newtonian force-fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):199– 208, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Einstein:1949:MPG

[20] Albert Einstein and Leopold Infeld. On the motion of particles in General Relativity Theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(1):209–241 + 1, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minakshisundaram:1949:SPE

[21] S. Minakshisundaram and Å. Pleijel. Some properties of the eigenfunctions of the Laplace-operator on Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):242–256, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Synge:1949:MTV

[22] J. L. Synge. On the motion of three vortices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1 (??):257–270, ???? 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weinstein:1949:SW

[23] Alexander Weinstein. On surface waves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1 (??):271–278, ???? 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Busemann:1949:AMI

[24] Herbert Busemann. Angular measure and integral curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):279–296, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chowla:1949:DRI

[25] S. D. Chowla and John Todd. The density of reducible integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):297–299, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taussky:1949:TLM

[26] Olga Taussky. On a theorem of Latimer and MacDuffee. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):300–302, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frame:1949:CRB

[27] J. S. Frame. Congruence relations between the traces of matrix powers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):303– 304, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorentz:1949:DTM

[28] G. G. Lorentz. Direct theorems on methods of summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):305–319, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minakshisundaram:1949:GEZ

[29] S. Minakshisundaram. A generalization of Epstein zeta functions. With a supplementary note by Hermann Weyl. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):320– 327, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mendelsohn:1949:ACF

[30] N. S. Mendelsohn. Applications of combinatorial formulae to generalizations of Wilson’s theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):328–336, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rado:1949:ATR

[31] R. Rado. Axiomatic treatment of rank in infinite sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):337–343, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pall:1949:RQF

[32] Gordon Pall. Representation by quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):344–364, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frucht:1949:GDT

[33] Robert Frucht. Graphs of degree three with a given abstract group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):365– 378, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1949:FLE

[34] G. F. D. Duff. Factorization ladders and eigenfunctions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):379–396, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mahler:1949:TLF

[35] K. Mahler. On a theorem of Liouville in fields of positive characteristic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):397– 400, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).